Given market demand Qd=50-p and market supply P=Qs+5
Find the market equilibrium price and quality
1
Expert's answer
2021-07-14T12:15:33-0400

If market supply is P = Qs + 5, then: Qs = P - 5

At market equilibrium: Qd = Qs

50 - P = P - 5

50 + 5 = P + P

2P = 55

P = 27.5

Pe = 27.5

Qe = 50 - (27.5)

Qe = 22.5

The price at market equilibrium is 27.5, the quantity at market equilibrium is 22.5